AAVSOnet status, 2016-01-13

Wed, 01/13/2016 - 13:06

BSM_HQ, BSM_NM, BSM_South and BSM_Berry are running nominally.  BSM_NM seems to have a very touchy focus and will drift out during the night.  We'll turn autofocus back on, once we solve a problem with that feature.   BSM_South seems to have some "banding" problems in its camera, not impacting photometry much, but visually noticable.  We're thinking about the problem.  BSM_Hamren's camera has been repaired and should be sent back to Mike Linnolt within a day or two.

OC61 had a sheared pin on the dome shutter, probably due to the high winds they have been experiencing.  That is now repaired, and no observing was lost since the two "down" nights were both poor.  Coker30 is working, but having poor weather.  TMO61 has a loose dome shutter bracket, at the top of the dome of course, and Jon Holtzman has not had it repaired yet.  It has been down since December 21.  Finally, SRO has been down since December 28, again due to dome problems.  The repaired control board has been shipped, and it should be back in action by the weekend.

A new focual reducer has been sent to OC61.  It will need to be adapted to the telescope, but is designed to improve the throughput of the echelle spectrograph.  With luck, we will be able to go 1-2 magnitudes fainter than its current limit.

We've been running standards at the BSM sites for the past month, so that we can redo the transformation coefficients.  Results of those efforts will be posted in a week or two.